OpenStudy (ujjwal):

A circular disk has 100 holes and it is rotating at the rate of 720 cycle/min..the frequency of emitted note will be a)1200 Hz b)70 Hz c)8400 Hz d)7.2 Hz

OpenStudy (vincent-lyon.fr):

I imagine a jet of compressed air is blown where the holes are? If not, I can't see how such a system could produce a sound wave.

OpenStudy (vincent-lyon.fr):

Work out how many times per second the air blows through a hole, and this is your frequency.

OpenStudy (ujjwal):

I guess the question says the air is blown 720 times per min, i.e 120 times per sec, through all 100 holes... I don't know how do solve this one..

OpenStudy (ujjwal):

@pastory i am sure you gave the answer by simply guessing.. But that's not in the list of options...

OpenStudy (jamesj):

@pastory, no that isn't right. Assuming the 100 holes are all at the edge, then in one minute, 720x100 = 72,000 holes are blown per minute or 72,000/60 = 12,000 per second.

OpenStudy (asnaseer):

720 cycles per minute, implies: 1 cycle in \(\frac{1}{720}\) minutes = \(\frac{1}{12}\)s therefore 360 degrees in \(\frac{1}{12}\)s therefore 1 degree in \(\frac{1}{12*360}\)s 100 holes would take up \(\frac{360}{100}\) degrees = 3.6 degrees therefore 3.6 degrees in \(\frac{3.6}{12*360}\)s = \(\frac{1}{1200}\)s and your frequency would be this fraction inverted, i.e. 1200Hz

OpenStudy (asnaseer):

JamesJ - I /think/ you forgot to cancel one zero

OpenStudy (asnaseer):

@ujjwal I used the fact that frequency \(f=\frac{1}{T}\)

OpenStudy (asnaseer):

sorry - in my explanation above I should have said "1 hole would take up..." instead of "100 holes would take up..."

OpenStudy (ujjwal):

100 holes would take up 360/100 degrees = 3.6 degrees I didn't get this one..

OpenStudy (asnaseer):

100 holes take up 360 degrees therefore each hole would be separated by \(\frac{360}{100}\) degrees

OpenStudy (ujjwal):

And that is frequency for rotation of each hole.. How does it end up as the frequency of note emitted?

OpenStudy (asnaseer):

think of the sinusoidal wave generated as each holes passes by

OpenStudy (asnaseer):

so you get a "peak" flow when the hole is in line with the air being blown, and it decays as the hole spins past

OpenStudy (asnaseer):

maybe someone else can give a better explanation?

OpenStudy (asnaseer):

imagine flicking your finger across your mouth as you blow out - you generate a sound wave whose frequency is equal to the frequency of your finger flicking

OpenStudy (ujjwal):

 which way is the disk rotating?